# Non-Stationary Covariance Functions for Spatial Data on Linear Networks

### v1

@@ 5 Discussion
This paper proposes non-stationary covariance functions for random fields defined on linear networks. The models exhibit local isotropy with respect to the resistance metric, allowing both the variance and correlation range to vary spatially. We demonstrate the effectiveness of a weighted local likelihood approach with compactly supported weights for estimating the proposed models, which provide a balance between computational efficiency and statistical accuracy.
Although constructive criticism has been raised regarding covariance models based on the resistance metric (Bolin et al., 2024), primarily because existing covariance families under this metric do not provide explicit control over the mean-square differentiability of the random field (i.e., there is no direct analogue of the Matérn covariance model under the resistance metric), we argue that the models introduced in this work offer considerable potential for practical applications. In the context of point process modeling, and in particular for LGCPs, local flexibility in variance and correlation range is essential for capturing heterogeneous aggregation structures, even when the smoothness level of the latent GRF is kept fixed (Dvořák et al., 2019). The proposed models are also applicable to geostatistical data, such as continuous measurements along stream networks. Since they do not allow control over the smoothness of the random field, their applicability is limited to phenomena exhibiting relatively rough variations across space. Extending the framework to include flexible smoothness represents a promising direction for future research.
Other strategies for modeling non-stationarity in Euclidean spaces, such as domain deformation or kernel convolutions of isotropic random fields, could potentially be adapted to linear networks (see Schmidt and Guttorp, 2020 and references therein). Investigating these approaches represents a promising direction for future research.
Another interesting avenue is the development of nonparametric and semiparametric estimation methods in the network setting, analogous to the approaches of Chang et al. (2010) and Kidd and Katzfuss (2022) in Euclidean spaces, as well as the design of fast geostatistical simulation algorithms by adapting the ideas in Alegría et al. (2026).

### v2

@@ 6 Discussion
This paper proposes non-stationary covariance functions for random fields defined on linear networks. The models exhibit local isotropy with respect to the resistance metric, allowing both the variance and correlation range to vary spatially. We provide a useful stochastic representation for these random fields and extend the covariance functions from the scalar to the matrix-valued setting, enabling the modeling of vector-valued random fields. We demonstrate the effectiveness of a weighted local likelihood approach with compactly supported weights for estimating the proposed covariance models, which provide a balance between computational efficiency and statistical accuracy.
Although constructive criticism has been raised regarding covariance functions based on the resistance metric (Bolin et al.,, 2024), primarily because existing covariance families under this metric do not provide explicit control over the mean-square differentiability of the random field (that is, there is no direct analogue of the Matérn family of covariance functions under the resistance metric), we argue that the models introduced in this work offer considerable potential for practical applications. In the context of point process modeling, and in particular for LGCPs, local flexibility in variance and correlation range is essential for capturing heterogeneous aggregation structures, even when the smoothness level of the latent GRF is kept fixed (Dvořák et al.,, 2019). The proposed covariance models are also applicable to geostatistical data, such as continuous measurements along stream networks (Ver Hoef and Peterson,, 2010). Since they do not allow control over the smoothness of the random field, their applicability is restricted to phenomena exhibiting relatively rough short-scale variations across space. Extending the framework to include flexible smoothness represents a promising direction for future research.
Other strategies for modeling non-stationarity in Euclidean spaces, such as domain deformation or different variants of kernel convolutions of isotropic random fields, could potentially be adapted to linear networks (see Schmidt and Guttorp,, 2020 and references therein). Investigating these approaches represents a promising direction for future research. Another interesting avenue is the development of nonparametric and semiparametric estimation methods in the network setting, analogous to the approaches of Chang et al., (2010) and Kidd and Katzfuss, (2022) in Euclidean spaces.
@@ Acknowledgement
Appendices
Appendix A Proof of Theorem 3.1
A direct calculation shows that, for all $w>0$,
$\frac{1}{\sqrt{\alpha(s,t)+\text{d}_{R}(s,t)w}}$ $=$ $\frac{1}{\sqrt{\alpha(s,t)}}\int_{\mathbb{R}}\exp\left\{-\frac{w\text{d}_{R}(s,t)y^{2}}{2\alpha(s,t)}\right\}\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{y^{2}}{2}\right\}\text{d}y,$ (A.1)
where the expectation in (A.1) is taken with respect to $Y\sim\mathcal{N}(0,1)$. Consequently, $K$ in Equation (3.1) can be written as
$K(s,t)=\frac{\sqrt{b(s)b(t)}}{\sqrt{\alpha(s,t)}}\int_{0}^{\infty}{\rm E}\left[\exp\left\{-\frac{w\text{d}_{R}(s,t)Y^{2}}{2\alpha(s,t)}\right\}\right]\text{d}F(w),\qquad s,t\in\mathcal{L}.$ (A.2)
Recall that $\phi(\cdot;\mu,\sigma^{2})$ denotes the probability density function of a normal distribution with mean $\mu$ and variance $\sigma^{2}$. Using a convolution argument, we obtain
$\int_{\mathbb{R}}\phi\left(z;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4w}\right)\phi\left(z;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4w}\right)\text{d}z=\frac{\sqrt{w}}{\sqrt{\pi\alpha(s,t)}}\exp\left\{-\frac{w\big[A(s)-A(t)\big]^{2}}{2\alpha(s,t)}\right\},$ (A.3)
for every $w>0$, with this expression conditioned on the values, $A(s)$ and $A(t)$, of the auxiliary GRF defined in Equation (2.1).
Notice that $A(s)-A(t)\sim\mathcal{N}(0,\mathrm{d}_{R}(s,t))$ and, consequently, $\big[A(s)-A(t)\big]/\sqrt{\mathrm{d}_{R}(s,t)}$ is equal in distribution to $Y\sim\mathcal{N}(0,1)$. Taking the expectation in (A.3) with respect to $(A(s),A(t))^{\top}$, we obtain
${\rm E}\left[\int_{\mathbb{R}}\phi\left(z;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4w}\right)\phi\left(z;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4w}\right)\text{d}z\right]=\frac{\sqrt{w}}{\sqrt{\pi\alpha(s,t)}}\,{\rm E}\left[\exp\left\{-\frac{w\text{d}_{R}(s,t)Y^{2}}{2\alpha(s,t)}\right\}\right],$ (A.4)
where the right-hand side of (A.4) simplifies to an expectation with respect to $Y\sim\mathcal{N}(0,1)$. Therefore, using Equation (A.4) and Fubini’s theorem, $K$ in (A.2) can be written as
$K(s,t)=\sqrt{\pi}\,{\rm E}\left[\int_{0}^{\infty}\int_{\mathbb{R}}\sqrt{b(s)b(t)}\,\phi\left(z;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4w}\right)\phi\left(z;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4w}\right)\text{d}z\,w^{-1/2}\,\text{d}F(w)\right].$ (A.5)
$\sum_{i=1}^{N}\sum_{j=1}^{N}r_{i}r_{j}K(s_{i},s_{j})= \sum_{i=1}^{N}\sum_{j=1}^{N}r_{i}r_{j}\sqrt{\pi}\,{\rm E}\bigg[\int_{0}^{\infty}\int_{\mathbb{R}}\sqrt{b(s_{i})b(s_{j})}\,\phi\bigg(z;\frac{A(s_{i})}{\sqrt{2}},\frac{a(s_{i})}{4w}\bigg)\\ \phi\bigg(z;\frac{A(s_{j})}{\sqrt{2}},\frac{a(s_{j})}{4w}\bigg)\text{d}z\,w^{-1/2}\,\text{d}F(w)\bigg],$
that is,
$\sum_{i=1}^{N}\sum_{j=1}^{N}r_{i}r_{j}K(s_{i},s_{j})=\sqrt{\pi}\,{\rm E}\left[\int_{0}^{\infty}\int_{\mathbb{R}}\left\{\sum_{i=1}^{N}r_{i}\sqrt{b(s_{i})}\,\phi\left(z;\frac{A(s_{i})}{\sqrt{2}},\frac{a(s_{i})}{4w}\right)\right\}^{2}\text{d}z\,w^{-1/2}\,\text{d}F(w)\right]\geq 0.$
Therefore, $K$ is positive semidefinite.
Appendix B Proof of Proposition 4.1
On the one hand, using the independence of the involved random variables and the fact that ${\rm E}[\epsilon]=0$, it is straightforward to see that the random field has zero mean. On the other hand, by independence and noting that ${\rm E}\big[\epsilon^{2}\big]=1$, the covariance function is given by
${\rm cov}\big[X(s),X(t)\big]$ $=$ ${\rm E}\big[\epsilon^{2}\big]{\rm E}\left[\frac{\sqrt{b(s)b(t)}}{g(Z)}\left(\frac{\pi}{W}\right)^{1/2}\phi\left(Z;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4W}\right)\phi\left(Z;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4W}\right)\right],$ (B.1)
$=$ ${\rm E}\left[\int_{0}^{\infty}\int_{\mathbb{R}}\frac{\sqrt{b(s)b(t)}}{g(z)}\left(\frac{\pi}{w}\right)^{1/2}\phi\left(z;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4w}\right)\phi\left(z;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4w}\right)g(z)\,\text{d}z\,\text{d}F(w)\right],$
$=$ ${\rm E}\left[\int_{0}^{\infty}\int_{\mathbb{R}}\sqrt{b(s)b(t)}\left(\frac{\pi}{w}\right)^{1/2}\phi\left(z;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4w}\right)\phi\left(z;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4w}\right)\,\text{d}z\,\text{d}F(w)\right],$
where the expectation in the last two equations is taken with respect to $(A(s),A(t))$. In Equation (B.1), we identify the expression obtained in Equation (A.5) in the proof of Theorem 3.1. Therefore, we recover the covariance function (3.1).
Appendix C Proof of Theorem 4.1
We first consider the simplified case where $\beta_{ij}=1$ for all $i,j=1,\ldots,q$. In this case, the function in Equation (4.2) reduces to $\tilde{K}_{ij}$ defined by
$\tilde{K}_{ij}(s,t)=\displaystyle\int_{0}^{\infty}\frac{\sqrt{b_{i}(s)b_{j}(t)}}{\sqrt{\alpha_{ij}(s,t)+\text{d}_{R}(s,t)w}}\text{d}F(w).$
Using the same arguments as in the proof of Theorem 3.1, we obtain
$\tilde{K}_{ij}(s,t)=\sqrt{\pi}\,{\rm E}\left[\int_{0}^{\infty}\int_{\mathbb{R}}\sqrt{b_{i}(s)b_{j}(t)}\,\phi\left(z;\frac{A(s)}{\sqrt{2}},\frac{a_{i}(s)}{4w}\right)\phi\left(z;\frac{A(t)}{\sqrt{2}},\frac{a_{j}(t)}{4w}\right)\text{d}z\,w^{-1/2}\,\text{d}F(w)\right].$ (C.1)
Note that this is a multivariate version of Equation (A.5). Let $N\in\mathbb{N}$, $\bm{r}_{1},\dots,\bm{r}_{N}\in\mathbb{R}^{q}$, and $s_{1},\dots,s_{N}\in\mathcal{L}$, where $\bm{r}_{k}=(r_{k1},\ldots,r_{kq})^{\top}$. Then,
$\sum_{k=1}^{N}\sum_{l=1}^{N}\bm{r}_{k}^{\top}\tilde{\bm{K}}(s_{k},s_{l})\bm{r}_{l}=\sum_{k=1}^{N}\sum_{l=1}^{N}\sum_{i=1}^{q}\sum_{j=1}^{q}r_{ki}r_{lj}\tilde{K}_{ij}(s_{k},s_{l})\\ =\sqrt{\pi}\,{\rm E}\left[\int_{0}^{\infty}\int_{\mathbb{R}}\left\{\sum_{k=1}^{N}\sum_{i=1}^{q}r_{ki}\sqrt{b_{i}(s_{k})}\,\phi\left(z;\frac{A(s_{k})}{\sqrt{2}},\frac{a_{i}(s_{k})}{4w}\right)\right\}^{2}\text{d}z\,w^{-1/2}\,\text{d}F(w)\right]\geq 0.$
In conclusion, the $Nq\times Nq$ matrix
$\begin{pmatrix}\tilde{\bm{K}}(s_{1},s_{1})&\cdots&\tilde{\bm{K}}(s_{1},s_{N})\\ \vdots&\ddots&\vdots\\ \tilde{\bm{K}}(s_{N},s_{1})&\cdots&\tilde{\bm{K}}(s_{N},s_{N})\end{pmatrix}$ (C.2)
is positive semidefinite, as expected.
For the general case, where the parameters $\beta_{ij}$ are not necessarily identically equal to one, we observe that
$\begin{pmatrix}{\bm{K}}(s_{1},s_{1})&\cdots&{\bm{K}}(s_{1},s_{N})\\ \vdots&\ddots&\vdots\\ {\bm{K}}(s_{N},s_{1})&\cdots&{\bm{K}}(s_{N},s_{N})\end{pmatrix}=\left[\bm{1}_{N}\bm{1}_{N}^{\top}\otimes\bm{\beta}\right]\circ\begin{pmatrix}\tilde{\bm{K}}(s_{1},s_{1})&\cdots&\tilde{\bm{K}}(s_{1},s_{N})\\ \vdots&\ddots&\vdots\\ \tilde{\bm{K}}(s_{N},s_{1})&\cdots&\tilde{\bm{K}}(s_{N},s_{N})\end{pmatrix}$, (C.3)
where $\circ$ and $\otimes$ denote the Hadamard and Kronecker products, respectively, and $\bm{1}_{N}$ is an $N$-dimensional vector of ones.
Since $\bm{\beta}$, $\bm{1}_{N}\bm{1}_{N}^{\top}$, and the matrix in (C.2) are positive semidefinite, and positive semidefiniteness is preserved under both Kronecker and Hadamard products, it follows that (C.3) is positive semidefinite.